Q\Determine the following ODEs are exact or not? Why?
(y^3)dx+3x^3y^2dy=0
1
Expert's answer
2020-10-26T19:58:30-0400

The given ODE is

"y^3 dx+3x^3y^2 dy = 0"



The necessary and sufficient condition that "M\space dx + N \space dy = 0" be exact is

"\frac{\delta M}{\delta y}=\frac{\delta N}{\delta x}"

"Here \space M= y^3 \space and \space N = 3x^3y^2"


"Now , \frac{\delta M}{\delta y}= 3y^2" ( partially differentiating M w.r.t 'y' keeping 'x' as constant)

"again, \frac{\delta N}{\delta x}= 9x^2y^2" ( partially differentiating N w.r.t 'x' keeping 'y' as constant )


As,

"\frac{\delta M}{\delta y} \neq\frac{\delta N}{\delta x}"

"\therefore" The given ODE is not exact.